Find the exact value without calculator:
(4cos30degrees - 6sin120degrees)^-2

Answer by Boreal(15235)

cos 30 is sqrt(3)/2
sin 120 is also sqrt (3)/2
(4sqrt(3)/2)-6 sqrt(3)/2)=-2 sqrt(3)/2
=-sqrt(3)
[-sqrt(3)]^(-2)=1/[-sqrt(3)]^2)=1/3